Use 9 unit multipliers to convert 10,000 cubic centimeters to cubic miles

Answer by CubeyThePenguin(3113)

1 cm^3 = 10^(-6) m^3
1 m^3 = 35.31 ft^3
1 ft^3 = 6.79 * 10^(-12) mi^3


10,000 cm^3 = 2.399 * 10^(-12) mi^3